# Calibrating Generative Models

### v1

@@ 4 Case-studies with diverse models, data, and constraints
We evaluate the capacity of CGM-reward and CGM-relax to solve practical calibration problems through three applications involving diverse model, data, and constraint types. Section˜4.1 calibrates a diffusion model (Lin et al., 2024a) and a masked language model (Hayes et al., 2025) of protein structure to more closely match statistics of natural proteins. Section˜4.2 calibrates a normalizing flow model (Zhai et al., 2025) of images to reduce class imbalances on the basis of LLM image-to-text annotations. Lastly, Section˜4.3 calibrates a small autoregressive LM to eliminate gender bias in generated children’s stories (Eldan & Li, 2023).
@@
Baselines. Only two prior works have proposed algorithms that intend to solve the calibration problem. Khalifa et al. (2021) propose a method for LLMs that we compare to in Section˜4.3. Second, Shen et al. (2024) propose a method for class-balancing in diffusion models. However, it assumes an existing probabilistic classifier and so is not applicable in our setting.
Compute cost. Each experiment is run on a single H100 GPU. We provide additional details regarding our experimental setup in Appendix˜E.
@@ 4.1 Calibrating protein design models to match statistics of natural proteins
Protein models Genie2 and ESM3 and their miscalibration. The two protein design models we consider are (1) Genie2 (Lin et al., 2024a), a 15M parameter equivariant diffusion model, and (2) ESM3-open (Hayes et al., 2025), a 1.4B parameter masked language model on tokenized representations of protein backbones. For each model, we generate protein backbones consisting of $100$ amino acids i.e. residues. Both Genie2 and ESM3-open suffer low diversity compared to natural protein domains in the CATH dataset (Sillitoe et al., 2021); specifically, they produce few generations with high beta-strand content (Figure˜3A). Beta strands, along with alpha helices and loops, constitute what is known as a protein’s secondary structure.
@@
Results. Performing calibration with CGM-relax yields a nearly fivefold improvement in the diversity of sampled protein structures for Genie2 and a twofold improvement for ESM3-open, as quantified by the symmetrized KL distance between the secondary structure distributions of the generative models and CATH domains (Figure˜3B). This improvement comes at the cost of an increased proportion of ‘design failures’, as defined in Section˜E.1. The ESM3-open base model generates a high proportion of design failures compared to Genie2 (consistent with Xiong et al. (2025), for example) and this fraction increases slightly upon calibration with CGM.
@@
The gains in secondary structure diversity achieved by CGM cannot be obtained by simply increasing the sampling noise of Genie2 or the sampling temperature of ESM3. In Figure˜3B, we show that increasing the sampling noise of Genie2 to $\sigma=1$ improves structure diversity, but at the cost of 5.3 times more design failures (failure rate 74%) than CGM. The same is true for ESM3 with increased sampling temperature $\tau=1$, which yields a $1.3$ times higher failure rate of $97\%$.
@@ 4.2 Calibrating Class Proportions in a Conditional Flow Model
We next demonstrate that CGM is capable of effectively calibrating state-of-the-art normalizing flow models. Normalizing flows generate samples $\boldsymbol{x}\in\mathbb{R}^{k}$ according to $\boldsymbol{x}=f_{\theta}^{-1}(\boldsymbol{\epsilon})$, where $\boldsymbol{\epsilon}\sim p_{\boldsymbol{\epsilon}}$ is a distribution from which sampling is tractable and $f_{\theta}(\boldsymbol{x})$ is a map that is invertible in $\boldsymbol{x}$ for each $\theta$ (Tabak & Vanden-Eijnden, 2010; Rezende & Mohamed, 2015). By the change-of-variable formula, the density of $\boldsymbol{x}$ is $p_{\boldsymbol{\epsilon}}(f_{\theta}(\boldsymbol{x}))|\text{det}(df_{\theta}(\boldsymbol{x})/d\boldsymbol{x})|$. This expression enables computation of exact likelihoods for maximum likelihood training and calibration.
For our calibration problem, we consider the 463M-parameter TarFlow model (Zhai et al., 2025), which parameterizes $f_{\theta}$ as an autoregressive vision transformer (Dosovitskiy et al., 2021) such that attention is performed over a sequence of image patches. We examine the model trained conditionally on the $256\times 256$ AFHQ dataset (Choi et al., 2020), which consists of images of animals faces belonging to one of three classes: $\{\texttt{cat, dog, wildlife}\}$. The wildlife class is further comprised of $\{\texttt{lion, tiger, fox, wolf, cheetah, leopard}\}$. We observe that, conditional on the wildlife class, approximately $36\%$ of generations from the TarFlow model are lions and very few ($7\%$ total) are foxes or wolves. We apply CGM to calibrate the conditional TarFlow model to generate samples containing animals from the wildlife class with equal proportions. For $\boldsymbol{h}$, we query GPT o5-mini to classify each image as containing one of the six animals or None.
Results. We find CGM-relax reduces miscalibration to the base TarFlow model with little visible degradation of sample realism (Figure˜4). CGM-relax $(\lambda{=}10^{-4})$ reduces the total variation distance of animal proportions, as classified by an image-to-text model, to the uniform distribution from $0.306$ to $0.108$. However, the Fréchet inception distance (FID) to real images in the AFHQ wildlife class is larger for the calibrated model than for the base model (21.9 vs. 15.9). Since this metric is sensitive to class proportions, we evaluate the calibrated model on the training dataset after balancing classes. The discrepancy in FID can be explained by two types of visual artifacts introduced by calibration: some images depict animals outside the wildlife class ($\sim 8\%$) and some “blend” multiple animals. Appendix Figure˜8 shows random samples from both models. The model fine-tuned with CGM-reward remains close to the base model but fails to reduce constraint violation.
@@ 4.3 Eliminating profession-specific gender imbalance in children’s stories
As a third example, we calibrate a language model that generates short children’s stories to remove gender bias. TinyStories-33M is an autoregressive transformer trained on children’s stories generated by GPT-3.5 and GPT4 (Eldan & Li, 2023). We find significant imbalances in prompt-conditional generations that introduce a character’s profession. For example, only 16% of stories beginning “Once upon a time there was a lawyer” feature a female lawyer, whereas 41% of U.S. attorneys were women in 2024 (American Bar Association, 2024).
@@
Gender parity as a calibration constraint and conditional calibration. We evaluate whether CGM can eliminate profession-specific gender imbalance in stories completed from the prompt “Once upon a time there was a <profession>” across eight professions that exhibit gender bias under the base model: doctor, lawyer, teacher, pilot, chef, scientist, nurse, and artist. In contrast to earlier experiments, this requires conditional calibration: for each profession $i$ with prompt $\text{prompt}_{i}$, we aim to find $\theta$ such that $\mathbb{E}_{p_{\theta}}[\boldsymbol{h}(\boldsymbol{x})~|~\text{prompt}_{i}]=0$, where $\boldsymbol{x}$ represents a completed story, and $\boldsymbol{h}(\boldsymbol{x})\in\{-1,0,1\}$ encodes the character’s gender (male, ambiguous, or female, respectively). Rather than fine-tuning a separate model for each profession, we amortize training costs by fine-tuning a single model with the sum of CGM losses for each condition.
@@
Results on explicitly calibrated professions. Both CGM-reward and CGM-relax reduce gender imbalance, as measured by the average absolute per-profession frequency difference (Figure˜5A). As expected, decreasing the regularization strength $\lambda$ improves constraint satisfaction at the cost of greater distance to the base-model, as measured by symmetrized KL. Notably, even the least-regularized model attains a low symmetrized KL of 1.7, which corresponds to an average token log-probability difference of <0.01<0.01 nats/token. Appendix E.4.4 provides example generations before and after fine-tuning showing no visible degradation in story quality.
@@
Compared to Khalifa et al. (2021), CGM-reward yields a small but statistically significant improvement in both miscalibration and distance to the base model. CGM-relax reduces gender imbalance by over five times more than Khalifa et al. (2021) but deviates further from the base-model.
@@
Figure 5: A: Gender imbalance and distance from base-model (symmetrized KL from pre-trained TinyStories-33M). B: Gender imbalance for professions included and heldout from calibration before and after CGM-relax ($\lambda=0.1$). Points below the diagonal were improved by CGM.
@@
Transference of calibration to heldout professions. We evaluate how conditional calibration affects the calibration of “held-out” professions not considered during fine-tuning. Such generalization could be particularly valuable in applications where it is impractical to foresee and explicitly calibrate for every possible prompt. To evaluate this, we consider six held-out professions: sheriff, judge, accountant, dancer, athlete, and baker. While CGM does not result in gender parity for the held-out professions, the imbalance is significantly reduced for some (Figure˜5B).
@@ 5 Conclusion
CGM-relax and CGM-reward provide practical approaches for calibrating generative models to satisfy distribution-level constraints. In applications to protein design, conditional image generation, and language modeling, CGM consistently reduces calibration error under hundreds of simultaneous constraints and in models with up to one billion parameters while preserving generation quality.
@@ 6 Acknowledgments
Appendix˜D: Simulation Experiments Additional Details
@@
    D.2: Initializing the Base Diffusion Model
    D.3: Experimental Details
@@
    E.1: Calibrating Genie2
    E.2: Calibrating ESM3-open
    E.3: Calibrating TarFlow
    E.4: Calibrating TinyStories-33M
@@ Appendix A Extended Discussion of Related Work
Empirically, the approximate solutions to the calibration problem (1) found by Khalifa et al. (2021) fall shorter of constraint satisfaction compared to CGM, particularly CGM-relax. Khalifa et al. (2021) achieves comparable, albeit slightly worse, performance to CGM-reward in the TinyStories gender rebalancing experiment (Section˜4.3), reducing miscalibration by roughly 85%. CGM-relax, on the other hand, reduces constraint violation up to 98%.
@@
Conditional generation (Dhariwal & Nichol, 2021; Ho & Salimans, 2021; Denker et al., 2024) can also be viewed through the lens of model calibration, where the calibration constraint is the indicator function of the set $C$ from which one would like to sample $\boldsymbol{h}(\boldsymbol{x})=\mathbbm{1}\{\boldsymbol{x}\in C\}$ and $\boldsymbol{h}^{\ast}$, the target proportion of samples that belong to $C$, approaches $1$. In this case the optimal variational parameter $\boldsymbol{\alpha}^{\ast}$ approach infinity, and the maximum entropy solution approaches $p_{\theta_{\text{base}}}(\boldsymbol{x})\mathbbm{1}\{\boldsymbol{x}\in C\}$.
@@
Calibration in prediction problems. Beyond generative modeling, calibration is a major topic in supervised machine-learning. In the context of classification the goal is to have, among a collection of predictions with a given class probability, the fraction of labels of that class in agreement with that prediction probability (Dawid, 1982). This can be obtained with post-hoc calibration procedures such as Platt scaling (Platt, 1999) or conformal methods (Shafer & Vovk, 2008) for more general prediction sets.
@@ B.2 Unbiased Gradient Estimates
The terms $(M-1)^{-1}\sum_{m^{\prime}\neq m}l_{m^{\prime}}$ and $(M-1)^{-1}\sum_{m^{\prime}\neq m}r_{m^{\prime}}$ in (LABEL:app:eq:gr-reward) are known as leave-one-out baselines (Kool et al., 2019) corresponding to sample $\boldsymbol{x}_{m}$. Including these terms adds to the score function gradient estimate a control variate, which is a term that has expectation zero under $p_{\theta}$ but is correlated with each individual term in the estimate (Lavenberg & Welch, 1981; Ranganath et al., 2014; Mohamed et al., 2020). Indeed, we observe that by independence of the samples $\{\boldsymbol{x}_{m}\}_{m=1}^{M}$, it holds that for each $m\neq m^{\prime}$,
@@
$\widehat{G}^{\text{relax}}=\nabla_{\theta}\widehat{\mathcal{L}}^{\text{viol}}\left(\left\{w(\theta,\theta^{\prime})\tilde{\boldsymbol{h}}_{m}\right\}_{m=1}^{M}\right)+\lambda\widehat{G}_{\text{KL}},\quad\boldsymbol{x}_{m}\overset{i.i.d.}{\sim}p_{\theta^{\prime}},\quad\theta^{\prime}=\texttt{stop-grad}(\theta).$
@@
From Section˜B.1, we know that $\widehat{G}_{\text{KL}}$ is unbiased for $\nabla_{\theta}\mathcal{L}^{\text{KL}}$, and so it only remains to verify that the second term is unbiased for $\nabla_{\theta}\mathcal{L}^{\text{viol}}=\nabla_{\theta}\|\mathbb{E}_{p_{\theta}}[\boldsymbol{h}]-\boldsymbol{h}^{\ast}\|^{2}$. To this end, by repeating the proof of Section˜B.1 (i.e. using the definition of the variance), it is straightforward to show

### v2

@@ 4 Case-studies with diverse models, data, and constraints
We evaluate the capacity of CGM-reward and CGM-relax to solve practical calibration problems through three applications involving diverse model, data, and constraint types. Section˜4.1 calibrates a diffusion model (Lin et al., 2024a) and a masked language model (Hayes et al., 2025) of protein structure to more closely match statistics of natural proteins. Section˜4.2 calibrates a normalizing flow model (Zhai et al., 2025) of images to reduce class imbalances on the basis of LLM image-to-text annotations. Lastly, Section˜4.3 calibrates a large language model to eliminate gender bias in generated children’s stories (Riviere et al., 2024).
@@
Baselines. Only two prior works have proposed algorithms that intend to solve the calibration problem. Khalifa et al. (2021) propose a method for LLMs that we compare to in Section˜4.3. Second, Shen et al. (2024) propose a method for class-balancing in diffusion models. However, their method assumes an existing probabilistic classifier and so is not applicable in our setting.
Compute cost. Each experiment is run on a single H100 GPU. Appendix˜E provides details on experimental setup.
@@ 4.1 Calibrating protein design models to match statistics of natural proteins
Protein models Genie2 and ESM3 and their miscalibration. The two protein design models we consider are (1) Genie2 (Lin et al., 2024a), a 15M parameter equivariant diffusion model, and (2) ESM3-open (Hayes et al., 2025), a 1.4B parameter masked language model on tokenized representations of protein backbones. For each model, we generate protein backbones consisting of $100$ amino acids i.e., residues. Both Genie2 and ESM3-open suffer low diversity compared to natural protein domains in the CATH dataset (Sillitoe et al., 2021); specifically, they produce few generations with high beta-strand content (Figure˜4A). Beta strands, along with alpha helices and loops, constitute what is known as a protein’s secondary structure.
@@
Results. Calibration with CGM-relax yields a nearly fivefold improvement in the diversity of sampled protein structures for Genie2 and a twofold improvement for ESM3-open, as quantified by the symmetrized KL distance between the secondary structure distributions of the generative models and CATH domains (Figure˜4B). This improvement comes at the cost of an increased proportion of ‘design failures’, as defined in Section˜E.2. The ESM3-open base model generates a high proportion of design failures compared to Genie2 (consistent with Xiong et al. (2025), for example) and this fraction increases slightly upon calibration with CGM.
@@
The gains in secondary structure diversity achieved by CGM cannot be obtained by simply increasing the sampling noise of Genie2 or the sampling temperature of ESM3. In Figure˜4B, we show that increasing the sampling noise of Genie2 to $\sigma=1$ improves structure diversity, but at the cost of a much higher failure rate (74% vs 14%) than CGM. The same is true for ESM3 with increased sampling temperature $\tau=1$, which yields a failure rate of $97\%$.
@@ 4.2 Calibrating Class Proportions in a Conditional Flow Model
We next demonstrate that CGM is capable of effectively calibrating state-of-the-art normalizing flow models. Normalizing flows generate samples $\boldsymbol{x}=f_{\theta}^{-1}(\boldsymbol{\epsilon})$, where $\boldsymbol{\epsilon}\sim p_{\boldsymbol{\epsilon}}$ is a distribution from which sampling is tractable and $f_{\theta}(\boldsymbol{x})$ is a map that is invertible in $\boldsymbol{x}$ for each $\theta$ (Tabak and Vanden-Eijnden, 2010; Rezende and Mohamed, 2015). By the change-of-variable formula, the density of $\boldsymbol{x}$ is $p_{\boldsymbol{\epsilon}}(f_{\theta}(\boldsymbol{x}))|\text{det}(df_{\theta}(\boldsymbol{x})/d\boldsymbol{x})|$. This expression enables computation of exact likelihoods for maximum likelihood training and calibration.
For our calibration problem, we consider the 463M-parameter TarFlow model (Zhai et al., 2025), which parameterizes $f_{\theta}$ as an autoregressive vision transformer (Dosovitskiy et al., 2021) that performs attention over a sequence of image patches. We examine the model trained conditionally on the $256\times 256$ AFHQ dataset (Choi et al., 2020), which consists of images of animals faces belonging to one of three classes: $\{\texttt{cat, dog, wildlife}\}$. The wildlife class is further comprised of $\{\texttt{lion, tiger, fox, wolf, cheetah, leopard}\}$. We observe that, conditional on the wildlife class, approximately $36\%$ of generations from the TarFlow model are lions and very few ($7\%$ total) are foxes or wolves. We apply CGM to calibrate the conditional TarFlow model to generate samples containing animals from the wildlife class with equal proportions. For $\boldsymbol{h}$, we query GPT o5-mini to classify each image as containing one of the six animals or None.
Results. We find CGM-relax reduces miscalibration to the base TarFlow model with little visible degradation of sample realism (Figure˜5). CGM-relax $(\lambda{=}10^{-4})$ reduces the total variation distance of animal proportions, as classified by an image-to-text model, to the uniform distribution from $0.306$ to $0.101$. However, the Fréchet inception distance (FID) to real images in the AFHQ wildlife class is larger for the calibrated model than for the base model ($21.0$ vs. $15.9$). Since this metric is sensitive to class proportions, we evaluate the calibrated model on the training dataset after balancing classes. The discrepancy in FID can be explained by two types of visual artifacts introduced by calibration: some images depict animals outside the wildlife class ($\sim 8\%$) and some “blend” multiple animals. Appendix Figure˜10 shows random samples from both models. The model fine-tuned with CGM-reward remains close to the base model but fails to reduce constraint violation.
@@ 4.3 Eliminating profession-specific gender imbalance in children’s stories
As a third example, we calibrate a large language model to generate short children’s stories with reduced gender bias using LoRA (Hu et al., 2022). Gemma-2-9B-IT is a nine billion parameter autoregressive transformer post-trained for instruction following Riviere et al. (2024). We find significant imbalances in prompt-conditional generations that introduce a character’s profession. For example, only 18% of stories beginning “Once upon a time there was a lawyer” feature a female lawyer, whereas 41% of U.S. attorneys were women in 2024 (American Bar Association, 2024).
Gender parity as a calibration constraint and conditional calibration. We evaluate whether CGM can eliminate profession-specific gender imbalance in childrens’ stories. We begin with a prompt adapted from Eldan and Li (2023):
Write a short story (3-5 paragraphs) which only uses very simple words that a 3 year old child would likely understand. ONLY write the story without any additional text. Try to use characters with roughly EQUAL PROBABILITIES male or female. The story begins: "Once upon a time there was a <profession> named
Compared to Eldan and Li (2023), our prompt differs in two ways. First, we prompt for stories about characters with eight professions that exhibit gender bias under the base model: doctor, lawyer, teacher, pilot, chef, scientist, nurse, and artist. And second, as a simpler strategy to reduce imbalance, we explicitly prompt for equal probabilities across genders; this reduces imbalance for some professions but generally does not eliminate the majority of miscalibration error (Section˜E.5.2).
@@
In contrast to earlier experiments, this requires conditional calibration: for each profession $i$ with prompt $\text{prompt}_{i}$, we aim to find $\theta$ such that $\mathbb{E}_{p_{\theta}}[\boldsymbol{h}(\boldsymbol{x})~|~\text{prompt}_{i}]=0$, where $\boldsymbol{x}$ represents a completed story, and $\boldsymbol{h}(\boldsymbol{x})\in\{-1,0,1\}$ encodes the character’s gender (male, ambiguous, or female, respectively). Rather than fine-tuning a separate model for each profession, we amortize training costs by fine-tuning a single model with the sum of CGM losses for each condition.
@@
Results on explicitly calibrated professions. Both CGM-reward and CGM-relax reduce gender imbalance, as measured by the average absolute per-profession frequency difference (Figure˜6A). As expected, decreasing the regularization strength $\lambda$ improves constraint satisfaction at the cost of greater distance to the base-model, as measured by symmetrized KL. Notably, even the least-regularized model attains a low symmetrized KL of <1.7{<}1.7, which corresponds to an average token log-probability difference of <0.01{<}0.01 nats/token. Appendix E.6.2 provides example generations before and after fine-tuning showing no visible degradation in story quality.
@@
Compared to Khalifa et al. (2021), CGM-reward yields a small but statistically significant improvement in miscalibration at the same distance to the base-model. CGM-relax reduces gender imbalance by over three times more than Khalifa et al. (2021) but deviates further from the base-model.
@@
Figure 6: A: Gender imbalance and distance from base-model (symmetrized KL from pre-trained Gemma-2-9B-IT). B: Gender imbalance for professions included and heldout from calibration before and after CGM-relax ($\lambda=0.01$). Points below the diagonal were improved by CGM.
@@
Transference of calibration to held-out professions. We evaluate how conditional calibration affects the calibration of “held-out” professions not considered during fine-tuning. Such generalization could be particularly valuable in applications where it is impractical to foresee and explicitly calibrate for every possible prompt. To evaluate this, we consider six held-out professions: sheriff, judge, accountant, dancer, athlete, and baker. Four of the six held-out professions are significantly improved, while two are made more imbalanced (Figure˜6B).
@@ 5 Conclusion
CGM-relax and CGM-reward provide practical approaches for calibrating generative models to satisfy distribution-level constraints. In applications to protein design, conditional image generation, and language modeling, CGM consistently reduces calibration error under hundreds of simultaneous constraints and in models with up to nine billion parameters while preserving generation quality.
@@ 6 Acknowledgments
Appendix˜D: Calibrating Continuous-time Diffusion Models
@@
    D.2: Solution to the Maximum Entropy Problem
    D.3: Exact Sampling from a Gaussian Mixture Model
@@
    E.1: Synthetic Data Experiments
    E.2: Calibrating Genie2
    E.3: Calibrating ESM3-open
    E.4: Calibrating TarFlow
    E.5: Calibrating Gemma-2-9B-IT
@@ Appendix A Extended Discussion of Related Work
Empirically, the approximate solutions to the calibration problem (1) found by Khalifa et al. (2021) fall shorter of constraint satisfaction compared to CGM, particularly CGM-relax. Khalifa et al. (2021) achieves comparable, albeit slightly worse, performance to CGM-reward in the Gemma 2 gender rebalancing experiment (Section˜4.3), reducing miscalibration by roughly 81%. CGM-relax, on the other hand, reduces constraint violation up to 94%.
@@
Conditional generation (Dhariwal and Nichol, 2021; Ho and Salimans, 2021; Denker et al., 2024) can also be viewed through the lens of model calibration, where the calibration constraint is the indicator function of the set $C$ from which one would like to sample $\boldsymbol{h}(\boldsymbol{x})=\mathbbm{1}\{\boldsymbol{x}\in C\}$ and $\boldsymbol{h}^{\ast}$, the target proportion of samples that belong to $C$, approaches $1$. In this case the optimal variational parameter $\boldsymbol{\alpha}^{\ast}$ approach infinity, and the maximum entropy solution approaches $p_{\theta_{\text{base}}}(\boldsymbol{x})\mathbbm{1}\{\boldsymbol{x}\in C\}$.
@@ B.2 Unbiased Gradient Estimates
The terms $(M-1)^{-1}\sum_{m^{\prime}\neq m}l_{m^{\prime}}$ and $(M-1)^{-1}\sum_{m^{\prime}\neq m}r_{m^{\prime}}$ in (LABEL:app:eq:gr-reward) are known as leave-one-out baselines (Kool et al., 2019) corresponding to sample $\boldsymbol{x}_{m}$. Including these terms adds to the score function gradient estimate a control variate, which is a term that has expectation zero under $p_{\theta}$ but is correlated with each individual term in the estimate (Lavenberg and Welch, 1981; Ranganath et al., 2014; Mohamed et al., 2020). Indeed, we observe that by independence of the samples $\{\boldsymbol{x}_{m}\}_{m=1}^{M}$, it holds that for each $m\neq m^{\prime}$,
@@
$\widehat{G}^{\text{relax}}=\nabla_{\theta}\widehat{\mathcal{L}}^{\text{viol}}\left(\left\{w_{m}(\theta,\theta^{\prime})\tilde{\boldsymbol{h}}_{m}\right\}_{m=1}^{M}\right)+\lambda\widehat{G}_{\text{KL}},\quad\boldsymbol{x}_{m}\overset{i.i.d.}{\sim}p_{\theta^{\prime}},\quad\theta^{\prime}=\texttt{stop-grad}(\theta).$
@@
From Section˜B.1, we know that $\widehat{G}_{\text{KL}}$ is unbiased for $\nabla_{\theta}\mathcal{L}^{\text{KL}}$, and so it only remains to verify that the second term is unbiased for $\nabla_{\theta}\mathcal{L}^{\text{viol}}=\nabla_{\theta}\|\mathbb{E}_{p_{\theta}}[\boldsymbol{h}]-\boldsymbol{h}^{\ast}\|^{2}$. To this end, by repeating the proof of Section˜B.1 (i.e., using the definition of the variance), it is straightforward to show